# Beware of CaBER: Filament thinning rheometry doesn’t give ‘the’ relaxation time of polymer solutions

### v3

@@ Abstract
The viscoelastic relaxation time of a polymer solution is often measured using Capillary Breakup Extensional Rheometry (CaBER) where a droplet is placed between two plates which are pulled apart to form a thinning filament. For a slow plate retraction protocol, required to avoid inertio-capillary oscillations for low-viscosity liquids, we show experimentally that the CaBER relaxation time $\tau_{e}$ inferred from the exponential thinning regime is in fact an apparent relaxation time that increases significantly when increasing the plate diameter and the droplet volume. Similarly, we observe that $\tau_{e}$ increases with the plate diameter for the classical step-strain plate separation protocol of a commercial (Haake) CaBER device and increases with the nozzle diameter for a Dripping-onto-Substrate (DoS) method. This dependence on the flow history before the formation of the viscoelastic filament is in contradiction with polymer models such as Oldroyd-B that predict a filament thinning rate $1/3\tau$ ($\tau$ being the model’s relaxation time) which is a material property independent of geometrical factors. We show that this is not due to artefacts such as solvent evaporation or polymer degradation and that it can only be rationalised by finite extensibility effects (FENE-P model) for a dilute polymer solution in a viscous solvent, but not for semi-dilute solutions in a low-viscosity solvent.
@@ 2.2 Slow stepwise plate separation CaBER protocol
Figure 1: (a) Time evolution of the minimum bridge / filament radius $h$ in the step-by-step plate separation protocol for the PEOaq solution for plate diameters $3.52R_{0}=$ 3.5 mm and a droplet volume $V^{*}=V/R_{0}^{3}\approx 2.4$. Inertio-capillary oscillations are visible after each step. Inset images correspond to a stable liquid bridge (left) and to a thinning filament (right) of the PEOaq solution with $72R_{0}=$ 7 mm and $V^{*}=2.4$. (b) $h(t)$ in log-lin for the PEOaq solution tested with plate diameters, $22R_{0}=$ 2, $3.5$, $5$ and $7$ mm, with $V^{*}\approx 2.4$. Inset images correspond to three times labelled 1 to 3 indicated on the $h(t)$ curve plus a fourth later time where $h$ is below our spatial resolution for $72R_{0}=$ 7 mm. The time reference $t_{1}$ marks the onset of the elastic regime.
@@ 3 Results
The apparent relaxation time $\tau_{e}$ measured with our slow stepwise plate separation CaBER protocol is plotted in figure 2(a) as a function of the initial bridge radius $h_{0}$ for all polymer solutions, plate diameters and droplet volumes, data points of the same colour corresponding to the same $R_{0}$ with different $V^{*}$. We observe that $\tau_{e}$ increases with both $R_{0}$ and $V^{*}$ and that all data points for a given solution collapse on a single curve when plotted against $h_{0}$, which is itself an increasing function of both $R_{0}$ and $V^{*}$. In other words, a given solution tested with two different ($R_{0},V^{*}$) sets but with the same $h_{0}$ yields the same $\tau_{e}$, as some examples show in figure 2(a). This suggests that $h_{0}$ is in fact the only relevant geometrical parameter of the problem. This is in agreement with the accepted idea that polymer deformations during capillary thinning are only influenced by the local extensional flow in the bridge / filament of maximum extension rate $\dot{\epsilon}=-2\dot{h}/h$ at its thinnest point, while the top and bottom end droplets act as passive liquid reservoirs, their size not directly influencing the pinch-off dynamics.
@@
Figure 2: Apparent relaxation time $\tau_{e}$ against $h_{0}$ for all three solutions for plate diameters $22R_{0}=$ 2, $3.5$, $5$ and $7$ mm (a) and for $22R_{0}=$ 2, $3.5$, $5$, $7$, $10$, $12.5$, $15$, $20$ and $25$ mm (b). In (a), data points of the same colour correspond to the same $R_{0}$ with different droplet volumes $V^{*}=V/R_{0}^{3}\approx 1.3$, $2.4$ and $3.2$. In (b), a single volume is used for each plate diameter ($V^{*}\approx 2.4$ for the smallest plates and $0.88$ for the largest plates). The inset images show stable liquid bridges for $22R_{0}=$ 2 mm (a) and $20$ mm (b). The linear fit is for the PEOaq solution for $h_{0}<2$ mm.
@@
As shown by the inset images in figure 2(a,b), the bottom end-drop becomes increasingly larger than the top one as $R_{0}$ increases since the Bond number $\mbox{{Bo}}=\rho gR_{0}^{2}/\gamma$ ranges between $0.16$ and $25$. However, the thinning dynamics is not driven by gravity since the “filament” Bond number $\mbox{{Bo}}_{f}=\rho gL_{f}h_{1}/\gamma$, comparing the typical capillary pressure $\gamma/h_{1}$ in the filament to the hydrostatic pressure $\rho gL_{f}$ over the filament length $L_{f}$, is only up to $0.1$ for the largest plates. This is also evident from the fact that filaments are not thicker at their base, see, e.g., the right inset image in figure 1(a).
@@ 4.2 Elasto-capillary balance with FENE-P
Figure 3: (a) Experimental $h(t)$ for the PEOaq solution tested with different plate diameters. The elastic regime ($t\geq t_{1}$) is fitted by equation 4 and 5 with relaxation time $\tau=\tau_{m}$ (from experiments, table 1), using $h_{1}$ and $A_{1}/L^{2}$ as fitting parameters. Fitting values of the parameter $A_{1}/L^{2}$ are given in the legend. (b) Experimental $h(t)$ for the PEOaq solution tested with a single plate diameter $102R_{0}=$ 10 mm and corresponding values of $A_{zz}(t)$ calculated from equation 2 with $\dot{\epsilon}=-2\dot{h}/h$ using the experimental values of $h$, with relaxation time $\tau=\tau_{m}$ (from experiments, table 1) where we vary the value of $L^{2}$ from $10$ to $10^{8}$ (FENE-P) and $L^{2}=\infty$ (Oldroyd-B), using $A_{zz}=1$ at $h=h_{0}$ as the initial condition.
@@ 4.4 Comparing fitting and calculated $A_{1}$ values
Figure 4: (a) Values of $A_{1}/L^{2}$ used as fitting parameters (light purple, see e.g. figure 3(a)), and values calculated from the FENE-P model (dark blue, see e.g. figure 3(b)) for $\tau=\tau_{m}$ and values of $L^{2}$ discussed in the text, against $h_{0}$ for all liquids and plate diameters. (b) Same values of $A_{1}$ calculated from the FENE-P model (dark blue), compared with $(h_{0}/h_{1})^{4}$ (light orange). All values of $A_{1}$ calculated from FENE-P are the same as in Oldroyd-B except for the PEOvisc solution at low $h_{0}$ where Oldroyd-B values are shown with empty blue symbols.
@@
In order to better understand this, we compare these calculated values of $A_{1}$ (dark blue) with their upper limit $(h_{0}/h_{1})^{4}$ (light orange) in figure 4(b). This upper limit corresponds to a relaxation time $\tau$ so large that polymer relaxation (right hand side of equation 2) is always negligible in the Newtonian regime ($t<t_{1}$), a case where equation 2 (with $\dot{\epsilon}=-2\dot{h}/h$) can be integrated as $A_{zz}h^{4}=h_{0}^{4}$. Differences in values of $(h_{0}/h_{1})^{4}$ among the three polymer solutions are due to differences in $h_{1}$ stemming from different elastic moduli $G$, as will be discussed in a separate paper where we show that the Oldroyd-B model gives $h_{1}\propto(GH^{4}/\gamma)^{1/3}$ where $H\to h_{0}$ for large relaxation times. We find in figure 4(b) that $A_{1}$ is very close to the $(h_{0}/h_{1})^{4}$ limit for the PEOaq and HPAM solutions at low $h_{0}$, meaning that polymer relaxation is indeed negligible, and that the ratio between the two increases as $h_{0}$ increases, meaning that relaxation becomes more important. This is consistent with the fact that the Deborah number $\mbox{{De}}=\tau_{m}/\tau_{R}$ decreases from $500$ ($\gg 1$, i.e., negligible relaxation) at the lowest $h_{0}$ to $5$ at the highest $h_{0}$, where $\tau_{R}=\sqrt{\rho h_{0}^{3}/\gamma}$ is the Rayleigh time scale relevant for the thinning dynamics of such low-viscosity liquids with Ohnesorge numbers $\mbox{{Oh}}=\eta_{0}/\sqrt{\rho\gamma h_{0}}$ which are up to $0.02$ for PEOaq and up to $0.1$ for HPAM (at the lowest $h_{0}$ values), i.e. $\mbox{{Oh}}\ll 1$. The reason why the calculated values of $A_{1}$ are fairly independent of $h_{0}$ for these two liquids, hence resulting in an impossible match with the fitting values of $A_{1}/L^{2}$ (light purple) in figure 4(a), is therefore because these calculated values of $A_{1}$ are close to their upper limit $(h_{0}/h_{1})^{4}$ which are fairly independent of $h_{0}$ themselves. The fact that $h_{1}$ if fairly proportional to $h_{0}$ in our experiments will be discussed in a separate paper.
@@
By contrast, the calculated values of $A_{1}$ for the PEOvisc solution are increasing as $h_{0}$ decreases, allowing for a good match with the fitting values of $A_{1}/L^{2}$ (light purple) in figure 4(a). This is because, unlike for the two other solutions, polymer relaxation (right hand side of equation 2) is not negligible in the Newtonian regime ($t<t_{1}$), as indicated by the greater difference between $A_{1}$ and the upper (relaxation-free) limit $(h_{0}/h_{1})^{4}$ for the PEOvisc solution in figure 4(b). This is due to the slower thinning dynamics in the Newtonian regime, caused by a larger shear viscosity with Ohnesorge numbers $\mbox{{Oh}}=\eta_{0}/\sqrt{\rho\gamma h_{0}}$ ranging between $0.6$ and $1.9$ in our range of $h_{0}$ values. Indeed, since all three solutions have comparable relaxation times, slower thinning dynamics mean that polymer relaxation is more important, i.e., $A_{1}<(h_{0}/h_{1})^{4}$. The reason why $A_{1}$ decreases as $h_{0}$ increases for the PEOvisc solution is because the ratio between $A_{1}$ and $(h_{0}/h_{1})^{4}$ increases with $h_{0}$. This is because relaxation is increasingly important as $h_{0}$ increases since the time scale of the Newtonian thinning dynamics, expected to scale as $\tau_{R}\propto h_{0}^{3/2}$ or as $\tau_{\mathrm{visc}}=\eta_{0}h_{0}/\gamma\propto h_{0}$ depending on Oh, increases with $h_{0}$, resulting in lower Deborah numbers (based on either $\tau_{R}$ or $\tau_{\mathrm{visc}}$).
@@ 4.5 Numerical simulations
Figure 5: $h(t)$ from experiments and simulations for the PEOvisc solution for different plate diameters. Simulations are performed for $22R_{0}=$ 2, $3.5$ and $5$ mm only, using the FENE-P model with $\tau=\tau_{m}$ and $L^{2}=2.0\times 10^{4}$. The elastic regime is fitted by equations 4 and 5 with $\tau=\tau_{m}$ where, like in figure 3(a), $A_{1}/L^{2}$ and $h_{1}$ are used as fitting parameters.
Hence, The FENE-P model can only explain the increase of the apparent relaxation time $\tau_{e}$ with $h_{0}$ for the PEOvisc solution which is the most dilute one and with the highest solvent viscosity out of our three solutions. This is checked further by numerical simulations of the axisymmetric problem (with gravity) using the full FENE-P constitutive equation. We use as fixed model parameters the values of $\eta_{s}$ and $\eta_{p}$ from the shear rheology, the high-$h_{0}$ limit $\tau=\tau_{m}$ for the relaxation time and the value of $L^{2}=2.0\times 10^{4}$ used in §4.4 (figure 4) for which the simplified analytical model in §4.2 and §4.3 could rationalise the apparent relaxation times $\tau_{e}$. The equations to be solved are the same as in Rubio et al [19] and the numerical methods are detailed in C. The initial condition is established by starting from a stable liquid bridge with a plate-to-plate distance $L_{p}$ just below the instability threshold value and slightly increasing $L_{p}$ to trigger the pinch-off.

### v4

# Beware of CaBER: Filament thinning rheometry does not always give ‘the’ relaxation time of polymer solutions

@@ Abstract
The viscoelastic relaxation time of a polymer solution is often measured using Capillary Breakup Extensional Rheometry (CaBER) where a droplet is placed between two plates which are pulled apart to form a thinning filament. For a slow plate retraction protocol, required to avoid inertio-capillary oscillations for low-viscosity liquids, we show experimentally that the CaBER relaxation time $\tau_{e}$ inferred from the exponential thinning regime is in fact an apparent relaxation time that may increase significantly when increasing the plate diameter and the droplet volume. Similarly, we observe that $\tau_{e}$ increases with the plate diameter for the classical step-strain plate separation protocol of a commercial (Haake) CaBER device and increases with the nozzle diameter for a Dripping-onto-Substrate (DoS) method. This dependence on the flow history before the formation of the viscoelastic filament is in contradiction with polymer models such as Oldroyd-B that predict a filament thinning rate $1/3\tau$ ($\tau$ being the model’s relaxation time) which is a material property independent of geometrical factors. We show that this is not due to artefacts such as solvent evaporation or polymer degradation and that it can only be rationalised by finite extensibility effects (FENE-P model) for a dilute polymer solution in a viscous solvent, but not for semi-dilute solutions in a low-viscosity solvent.
@@ 2.2 Slow stepwise plate separation CaBER protocol
Figure 1: (a) Time evolution of the minimum bridge / filament radius $h$ in the slow stepwise plate separation protocol for the PEOaq solution for plate diameters $3.52R_{0}=$ 3.5 mm and a droplet volume $V^{*}=V/R_{0}^{3}\approx 2.4$. Inertio-capillary oscillations are visible after each step. Inset images correspond to a stable liquid bridge (left) and to a thinning filament (right) of the PEOaq solution with $72R_{0}=$ 7 mm and $V^{*}\approx 2.4$. (b) $h(t)$ in log-lin for the PEOaq solution tested with plate diameters, $22R_{0}=$ 2, $3.5$, $5$ and $7$ mm, with $V^{*}\approx 2.4$. Inset images correspond to three times labelled 1 to 3 indicated on the $h(t)$ curve plus a fourth later time where $h$ is below our spatial resolution for $72R_{0}=$ 7 mm. The time reference $t_{1}$ marks the onset of the elastic regime.
@@ 3 Results
The apparent relaxation time $\tau_{e}$ measured with our slow stepwise plate separation CaBER protocol is plotted in figure 2(a) as a function of the initial bridge radius $h_{0}$ for various plate diameters and droplet volumes for all three polymer solutions, data points of the same colour corresponding to the same $R_{0}$ with different $V^{*}$. We observe that $\tau_{e}$ increases with both $R_{0}$ and $V^{*}$ and that all data points for a given solution collapse on a single curve when plotted against $h_{0}$, which is itself an increasing function of both $R_{0}$ and $V^{*}$. In other words, a given solution tested with two different ($R_{0},V^{*}$) sets but with the same $h_{0}$ yields the same $\tau_{e}$, as some examples show in figure 2(a). This suggests that $h_{0}$ is in fact the only relevant geometrical parameter of the problem. This is in agreement with the accepted idea that polymer deformations during capillary thinning are only influenced by the local extensional flow in the bridge / filament of maximum extension rate $\dot{\epsilon}=-2\dot{h}/h$ at its thinnest point, while the top and bottom end droplets act as passive liquid reservoirs, their size not directly influencing the pinch-off dynamics.
Figure 2: Apparent relaxation time $\tau_{e}$ against $h_{0}$ for all three solutions for plate diameters $22R_{0}=$ 2, $3.5$, $5$ and $7$ mm (a) and for $22R_{0}=$ 2, $3.5$, $5$, $7$, $10$, $12.5$, $15$, $20$ and $25$ mm (b). In (a), data points of the same colour correspond to the same $R_{0}$ with different droplet volumes $V^{*}=V/R_{0}^{3}\approx 1.3$, $2.4$ and $3.2$ (only $V^{*}\approx 2.4$ for the HPAM solution). In (b), a single volume is used for each plate diameter ($V^{*}\approx 2.4$ for the smallest plates and $0.88$ for the largest plates). The inset images show stable liquid bridges for $22R_{0}=$ 2 mm (a) and $20$ mm (b). The linear fit is for the PEOaq solution for $h_{0}<2$ mm.
@@
As shown by the inset images in figure 2(a,b), the bottom end-drop becomes increasingly larger than the top one as $R_{0}$ increases since the Bond number $\mbox{{Bo}}=\rho gR_{0}^{2}/\gamma$ ranges between $0.16$ and $25$. However, the thinning dynamics is not driven by gravity since the “filament” Bond number $\mbox{{Bo}}_{f}=\rho gL_{f}h_{1}/\gamma$, comparing the typical capillary pressure $\gamma/h_{1}$ in the filament to the hydrostatic pressure $\rho gL_{f}$ over the filament length $L_{f}$, is only up to $0.1$ for the largest plates. This is also evident from the fact that filaments are not thicker at their base, see, e.g., the right inset image in figure 1(a). This is consistent with the fact that, as we discuss in our longer companion paper [15], the bridge thinning dynamics is well captured by the classical inertio-capillary and visco-capillary (gravity-free) self-similar laws (close to the transition to the elastic regime) for the PEOaq and PEOvisco solutions, respectively, for plate diameters where $\tau_{e}$ depends on $h_{0}$ (hence suggesting that the $h_{0}$-dependence of $\tau_{e}$ is not caused by gravitational drainage). These apparent extensional relaxation times $\tau_{e}$ are compared with relaxation times inferred from shear rheology in our companion paper [15].
@@ 4.2 Elasto-capillary balance with FENE-P
Figure 3: (a) Experimental $h(t)$ for the PEOaq solution tested with different plate diameters. The elastic regime ($t\geq t_{1}$) is fitted by equation 4 and 5 with relaxation time $\tau=\tau_{m}$ (from experiments, table 1), using $h_{1}$ and $A_{1}/L^{2}$ as fitting parameters. Fitting values of the parameter $A_{1}/L^{2}$ are given in the legend. (b) Experimental $h(t)$ for the PEOaq solution tested with a single plate diameter $102R_{0}=$ 10 mm and corresponding values of $A_{zz}(t)$ calculated from equation 2 with $\dot{\epsilon}=-2\dot{h}/h$ using the experimental values of $h$, with relaxation time $\tau=\tau_{m}$ (from experiments, table 1) where we vary the value of $L^{2}$ from $10$ to $10^{8}$ (FENE-P) and $L^{2}=\infty$ (Oldroyd-B), using $A_{zz}=1$ at $h=h_{0}$ as the initial condition. The Ohnesorge numbers $\mbox{{Oh}}=\eta_{0}/\sqrt{\rho\gamma h_{0}}$ ranges between $0.02$ and $0.007$ for the PEOaq solution in our range of $h_{0}$ values which corresponds to plates diameters $2R_{0}$ between $2$ and $25$ mm.
@@ 4.4 Comparing fitting and calculated $A_{1}$ values
Figure 4: (a) Values of $A_{1}/L^{2}$ used as fitting parameters (light purple, see e.g. figure 3(a)), and values calculated from the FENE-P model (dark blue, see e.g. figure 3(b)) for $\tau=\tau_{m}$ and values of $L^{2}$ discussed in the text, against $h_{0}$ for all liquids and plate diameters. (b) Same values of $A_{1}$ calculated from the FENE-P model (dark blue), compared with $(h_{0}/h_{1})^{4}$ (light orange). All values of $A_{1}$ calculated from FENE-P are the same as in Oldroyd-B except for the PEOvisc solution at low $h_{0}$ where Oldroyd-B values are shown with empty blue symbols. The Ohnesorge numbers $\mbox{{Oh}}=\eta_{0}/\sqrt{\rho\gamma h_{0}}$ ranges between $0.02$ and $0.007$ for the PEOaq solution, between $0.1$ and $0.03$ for the HPAM solution and between $2$ and $0.6$ for the PEOvisc solution in our range of $h_{0}$ values which corresponds to plates diameters $2R_{0}$ between $2$ and $25$ mm.
@@
In order to better understand this, we compare these calculated values of $A_{1}$ (dark blue) with their upper limit $(h_{0}/h_{1})^{4}$ (light orange) in figure 4(b). This upper limit corresponds to a relaxation time $\tau$ so large that polymer relaxation (right hand side of equation 2) is always negligible in the Newtonian regime ($t<t_{1}$), a case where equation 2 (with $\dot{\epsilon}=-2\dot{h}/h$) can be integrated as $A_{zz}h^{4}=h_{0}^{4}$. Differences in values of $(h_{0}/h_{1})^{4}$ among the three polymer solutions are due to differences in $h_{1}$ stemming from different elastic moduli $G$, as we discuss in our longer companion paper [15] where we show that the Oldroyd-B model gives $h_{1}\propto(GH^{4}/\gamma)^{1/3}$ where $H\to h_{0}$ for large relaxation times. We find in figure 4(b) that $A_{1}$ is very close to the $(h_{0}/h_{1})^{4}$ limit for the PEOaq and HPAM solutions at low $h_{0}$, meaning that polymer relaxation is indeed negligible, and that the ratio between the two increases as $h_{0}$ increases, meaning that relaxation becomes more important. This is consistent with the fact that the Deborah number $\mbox{{De}}=\tau_{m}/\tau_{R}$ decreases from $500$ ($\gg 1$, i.e., negligible relaxation) at the lowest $h_{0}$ to $5$ at the highest $h_{0}$, where $\tau_{R}=\sqrt{\rho h_{0}^{3}/\gamma}$ is the Rayleigh time scale relevant for the thinning dynamics of such low-viscosity liquids, with Ohnesorge numbers $\mbox{{Oh}}=\eta_{0}/\sqrt{\rho\gamma h_{0}}$ which are up to $0.02$ for PEOaq and up to $0.1$ for HPAM (at the lowest $h_{0}$ values), i.e. $\mbox{{Oh}}\ll 1$. The reason why the calculated values of $A_{1}$ are fairly independent of $h_{0}$ for these two liquids, hence resulting in an impossible match with the fitting values of $A_{1}/L^{2}$ (light purple) in figure 4(a), is therefore because these calculated values of $A_{1}$ are close to their upper limit $(h_{0}/h_{1})^{4}$ which are fairly independent of $h_{0}$ themselves (at least for sufficiently low $h_{0}$, see figure 4(b)). Note that the scaling $h_{1}\propto h_{0}$ implied by the constant values of $(h_{0}/h_{1})^{4}$ is discussed in our longer companion paper [15].
@@
By contrast, the calculated values of $A_{1}$ for the PEOvisc solution are increasing as $h_{0}$ decreases, allowing for a good match with the fitting values of $A_{1}/L^{2}$ (light purple) in figure 4(a). This is because, unlike for the two other solutions, polymer relaxation (right hand side of equation 2) is not negligible in the Newtonian regime ($t<t_{1}$), as indicated by the greater difference between $A_{1}$ and the upper (relaxation-free) limit $(h_{0}/h_{1})^{4}$ for the PEOvisc solution in figure 4(b). This is due to the slower thinning dynamics in the Newtonian regime, caused by a larger shear viscosity with Ohnesorge numbers $\mbox{{Oh}}=\eta_{0}/\sqrt{\rho\gamma h_{0}}$ ranging between $0.6$ and $2$ in our range of $h_{0}$ values. Indeed, since all three solutions have comparable relaxation times, slower thinning dynamics mean that polymer relaxation is more important, i.e., $A_{1}<(h_{0}/h_{1})^{4}$. The reason why $A_{1}$ decreases as $h_{0}$ increases for the PEOvisc solution is because the ratio between $A_{1}$ and $(h_{0}/h_{1})^{4}$ increases with $h_{0}$. This is because relaxation is increasingly important as $h_{0}$ increases since the time scale of the Newtonian thinning dynamics, expected to scale as $\tau_{R}\propto h_{0}^{3/2}$ or as $\tau_{\mathrm{visc}}=\eta_{0}h_{0}/\gamma\propto h_{0}$ depending on Oh, increases with $h_{0}$, resulting in lower Deborah numbers (based on either $\tau_{R}$ or $\tau_{\mathrm{visc}}$).
@@ 4.5 Numerical simulations
Figure 5: $h(t)$ from experiments and simulations for the PEOvisc solution for different plate diameters. Simulations are performed for $22R_{0}=$ 2, $3.5$ and $5$ mm only, using the FENE-P model with $\tau=\tau_{m}$ and $L^{2}=2.0\times 10^{4}$. The elastic regime is fitted by equations 4 and 5 with $\tau=\tau_{m}$ where, like in figure 3(a), $A_{1}/L^{2}$ and $h_{1}$ are used as fitting parameters. The Ohnesorge numbers $\mbox{{Oh}}=\eta_{0}/\sqrt{\rho\gamma h_{0}}$ ranges between $2$ and $0.6$ for the PEOvisc solution in our range of $h_{0}$ values which corresponds to plates diameters $2R_{0}$ between $2$ and $25$ mm.
Hence, The FENE-P model can only explain the increase of the apparent relaxation time $\tau_{e}$ with $h_{0}$ for the PEOvisc solution which is the most dilute one and with the highest solvent viscosity out of our three solutions. This is checked further by numerical simulations of the axisymmetric problem (with gravity) using the full FENE-P constitutive equation. We use as fixed model parameters the values of $\eta_{s}$ and $\eta_{p}$ from the shear rheology, the high-$h_{0}$ limit $\tau=\tau_{m}$ for the relaxation time and the value of $L^{2}=2.0\times 10^{4}$ used in §4.4 (figure 4(a)) for which the simplified analytical model in §4.2 and §4.3 could rationalise the apparent relaxation times $\tau_{e}$. The equations to be solved are the same as in Rubio et al [20] and the numerical methods are detailed in C. The initial condition is established by starting from a stable liquid bridge with a plate-to-plate distance $L_{p}$ just below the instability threshold value and slightly increasing $L_{p}$ to trigger the pinch-off.